OpenStudy (anonymous):

evaluate the integral by making the given substitution. int[(5*sin*sqrt(x))/sqrt(x)] , u=sqrt(x)

OpenStudy (anonymous):

forgot the dx! int[(5*sin*sqrt(x))/sqrt(x)] *dx , u=sqrt(x)

OpenStudy (anonymous):

\[5\int \frac{\sin(\sqrt{x})}{\sqrt{x}}dx\]

OpenStudy (anonymous):

\[u=\sqrt{x}\] \[du=\frac{1}{2\sqrt{x}}dx\] \[2du=\frac{dx}{\sqrt{x}}\] \[10\int \sin(u)du\]

OpenStudy (anonymous):

\[-10\cos(u)\] \[-10\cos(\sqrt{x})+C\]

OpenStudy (anonymous):

check is to take the derivative and see that you get the integrand

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

yw